ROI and Investment Decisions

Jarriot, Inc., presented two years of data for its Furniture Division and its Houseware Division.

Furniture Division:
            Year 1 Year 2
Sales $35,600,000 $37,600,000
Operating income 1,400,000 1,520,000
Average operating assets 4,200,000 4,200,000
Houseware Division:
            Year 1 Year 2
Sales $12,000,000 $12,500,000
Operating income 600,000 530,000
Average operating assets 5,550,000 5,550,000

At the end of Year 2, the manager of the Houseware Division is concerned about the division’s performance. As a result, he is considering the opportunity to invest in two independent projects. The first is called the Espresso-Pro; it is an in-home espresso maker that can brew regular coffee as well as make espresso and latte drinks. While the market for espresso drinkers is small initially, he believes this market can grow, especially around gift-giving occasions. The second is the Mini-Prep appliance that can be used to do small chopping and dicing chores that do not require a full-sized food processor. Without the investments, the division expects that Year 2 data will remain unchanged. The expected operating incomes and the outlay required for each investment are as follows:

Espresso-Pro Mini-Prep
Operating income $28,000 $15,500
Outlay 210,000 160,000

Jarriot’s corporate headquarters has made available up to $540,000 of capital for this division. Any funds not invested by the division will be retained by headquarters and invested to earn the company’s minimum required rate of return, 10 percent.

Required:

Round your answers to four decimal places before converting to a percentage. For example, .06349 would be rounded to .0635 and entered as "6.35" percent.

1. Compute the ROI for each investment.

Espresso-Pro ROI %
Mini-Prep ROI %

2. Compute the divisional ROI for each of the following four alternatives:

a. The Espresso-Pro is added.
%

b. The Mini-Prep is added.
%

c. Both investments are added.
%

d. Neither investment is made; the status quo is maintained.
%

Assuming that divisional managers are evaluated and rewarded on the basis of ROI performance, which alternative do you think the divisional manager will choose?
Answer #1
Part-1
Espresso-Pro ROI (28000/210000)= 13.33%
Mini-prep ROI (15500/160000)=9.69%
Part-2
a Espresso pro is added
(28000+530000)/(210000+5550000)=9.69%
b Mini prep is added
(15500+530000)/(160000+5550000)=9.55%
c Both are added
(15500+28000+530000)/(210000+160000+5550000)=9.69%
d Status quo (530000/5550000)=9.54%
ROI = Operating income/Operating assets
Espresso-Pro should be chosen as ROI is highest